OpenStudy (anonymous):

find the exact value : sin (-620 degrees)

OpenStudy (dumbcow):

sin(-620) = sin(720-620) = sin(100)

OpenStudy (anonymous):

why does 720 subtract from 620

OpenStudy (mertsj):

Because 720 is two complete revolutions.

OpenStudy (mertsj):

And you want to find the reference angle

OpenStudy (anonymous):

and 100 ends up being the exact value?

OpenStudy (mertsj):

sin(100)=sin(80)= exact value.

OpenStudy (mertsj):

Are you sure you didn't make a typo?

OpenStudy (anonymous):

wait a minute, yes its 630 degrees.

OpenStudy (mertsj):

Is it negative 630?

OpenStudy (anonymous):

yes still negative it says follow the even- odd properties

OpenStudy (anonymous):

it was an exam question i got wrong, and i put -1 as an answer because i thought it was sin 90 degrees..i used my calculator to answer

OpenStudy (mertsj):

I don't know what you mean by the even-odd properties but ...Start with -630 add 360 so now you are back at your starting point and still have -270 degrees to go. That brings you to 90 degrees and the sin is 1

OpenStudy (mertsj):